A motorboat can maintain a constant speed of 38 miles per hour relative to the water. The boat makes a trip upstream to a certain point in 55 minutes; the return trip take 40 minutes. what is the speed of the current?
Answer by checkley79(3341)

D=RT
D=(38-C)55/60
D=(44+C)40/60
55(38-C/60=40(44+C)/60 CANCEL OUT THE DENOMINATORS.
55(38-C)=40(44+C)
2090-55C=1760+40C
-55C-40C=1760-2090
-95C=-330
C=-330/-95
C=3.47 MPH. FOR THE CURRENT.
PROOF:
55(38-3.47)=40(44+3.47)
55*34.53=40*47.47
1899=1899